# Connection-Aware P2P Trading: Simultaneous Trading and Peer Selection

### v1

@@ Abstract
Index Terms:
P2P Energy Trading, Distributed Energy Resource, Trading Connection, Self-Selection Rights, Communication Asynchrony.
I Introduction
@@
Recent research highlights the importance of P2P trading in future distribution networks. Part of this research focused on a more detailed modeling of prosumers, considering their strategic behaviors [8] and bounded rationality [9]. Others incorporated a more detailed network model into trading problems, employing sensitivity factors [10] and DistFlow [11]. These models are valuable references for P2P trading analysis. Meanwhile, the implementation and development of solutions to trading models is also of significant importance. Various information exchange structures are designed to facilitate P2P trading, including leader-follower [12], bilateral [13], multi-layer [14], and grand coalition schemes [15]. This paper adopts a decentralized structure for its privacy protection and independence from a coordination entity [16]. Notably, most of the trading algorithms mentioned above, regardless of the information structure, neglect connection limits and require strict synchronization between trading peers or between the prosumer and the coordinator.
@@
Addressing connection issues in P2P trading is critical. In coalition-based P2P trading, some studies [17, 18] illustrated that the trading connections among prosumers are flexible and can be adjusted according to prosumers’ preferences. One study proposes a graph coalition formation game [17], while another suggests peer matching [18]. The same problem is true for non-coalition P2P trading, yet solutions to this issue are scarce. A few studies have investigated the non-coalition P2P trading connection problem. Ref.[19] offered an overview of information-related issues in P2P trading, while Ref.[20] conducted extensive numerical experiments on this topic. Some works [21, 22] adopted an even-triggered version of the alternating direction method of multipliers (ADMM) algorithm to address connection issues, named Communication-Sensored Consensus ADMM (COCA) [23]. In COCA, the connection peers are activated based on a diminishing threshold. Like all event-trigger algorithms, determining this threshold in real-world scenarios is challenging. Besides, the diminishing threshold should converge to zero, potentially activating almost all links over time. For asynchrony, Ref.[24, 25] pioneered the use of ADMM to counter asynchronous P2P trading. Nevertheless, Ref.[24] lacked a theoretical convergence guarantee for the constrained P2P trading problem. Ref.[25] introduced complex consensus processes during trading, exacerbating connection issues. Crucially, these works all assumed random connection problems, which ignored prosumers’ willingness and overlooked their self-selection rights.
@@
The remainder of this paper is organized as follows: Section II outlines the P2P trading framework, formulates the problem of computing the P2P trading equilibrium, and discusses the standard primal-dual solution algorithm. Section III introduces two variants of the asynchronous P2P trading algorithm - the edge-based and node-based algorithms - and elaborates on the smart selection strategies. Section IV presents a convergence proof for both algorithms and demonstrates the efficacy of the smart selection strategies in accelerating convergence. Section V conducts case studies validating the effectiveness of the proposed methods. Section VI offers conclusions and suggests avenues for future research.
@@
II Standard P2P Trading Model and Algorithm
II-A Trading Setting and Competitive Equilibrium Computation
The system comprises prosumers, denoted as $i=1,...,I$. Each prosumer $i$ possesses DERs, including load equipment, rooftop photovoltaic panels, electric vehicles, and energy storage systems. These are integrated with an energy management system to facilitate monitoring, P2P trading, and billing. In the context of P2P trading, ponsumers $i$ are represented as nodes $\mathcal{N}$, and potential trading connections are symbolized by $\mathcal{E}$, following the conventions of graph theory. This study considers the situation where P2P trading is carried out in the day-ahead market, where $\tau=1,...,\bar{\tau}$ periods are considered. For prosumer $i$, its individual decision variables can include load usage power $\boldsymbol{p}_{i}^{\mathrm{L}}=\mathrm{col}(p_{i,\tau}^{\mathrm{L}})$, storage/vehicle charging power $\boldsymbol{p}_{i}^{\mathrm{CH}}=\mathrm{col}(p_{i,\tau}^{\mathrm{CH}})$, storage/vehicle discharging power $\boldsymbol{p}_{i}^{\mathrm{DIS}}=\mathrm{col}(p_{i,\tau}^{\mathrm{DIS}})$, storage/vehicle state of charge $\boldsymbol{s}_{i}=\mathrm{col}(s_{i,\tau})$, and exchanging power with the main grid $\boldsymbol{p}_{i}^{\mathrm{EX}}=\mathrm{col}(p_{i,\tau}^{\mathrm{EX}})$. For brevity, they are denoted in the concatenated form $\boldsymbol{x}_{i}$. $\boldsymbol{x}_{i}=\mathrm{col}(\boldsymbol{p}_{i}^{\mathrm{L}},\boldsymbol{p}_{i}^{\mathrm{CH}},\boldsymbol{p}_{i}^{\mathrm{DIS}},\boldsymbol{s}_{i},\boldsymbol{p}_{i}^{\mathrm{EX}})$. The information is private to prosumer $i$. P2P trading transactions between prosumer $i$ and $j$ are denoted as $\boldsymbol{t}_{i,j}=\mathrm{col}({t}_{i,j,\tau})$ for prosumer $i$ and $\boldsymbol{t}_{j,i}=\mathrm{col}({t}_{j,i,\tau})$ for prosumer $j$, which should only be known by $i$ and $j$. All possible P2P trading transactions of prosumer $i$ are denoted as $\boldsymbol{t}_{i}=\mathrm{col}(\boldsymbol{t}_{i,j})_{j\in\mathcal{N}_{i}}$.
It is widely recognized that in a system with a large number of prosumers, the trading outcome will converge to a competitive equilibrium or Walrasian equilibrium [26], which is the optimal solution to the social cost minimization problem, represented as:
@@ II-A Trading Setting and Competitive Equilibrium Computation
where the objective function (1) is to minimize the sum of individual costs $J_{i}\left(\boldsymbol{x}_{i},\boldsymbol{t}_{i}\right)$, which further depends on $i$’s decision variables $\boldsymbol{x}_{i}$ and P2P transactions $\boldsymbol{t}_{i}$. The cost typically includes negative load usage utility, storage age cost, network fee, tax and preferences for P2P trading, etc. Constraint (2) is the operation constraint for prosumer $i$, denoted by $\Omega_{i}$. The constraints can include load shifting constraints, energy storage charging status constraints, and power output constraints, etc. All the related information belongs to prosumer $i$’s privacy. Constraint (3) is the P2P trading power balance constraint: the power exported from $i$ to $j$ ($\boldsymbol{t}_{i,j}$) should be balanced by power from $j$ from $i$ ($\boldsymbol{t}_{j,i}$). The corresponding dual variable is $\boldsymbol{\lambda}_{i,j}$, represented by $\boldsymbol{\lambda}_{i,j}=\mathrm{col}(\lambda_{i,j,\tau})$ in different trading periods.
@@
II-B Primal-dual Algorithm for P2P trading
@@ 3       Step 2: Each prosumer receives new proposals and updates dual prices
Moreover, the primal-dual algorithm can be interpreted as prosumers engaging in a potential game, where the potential function aligns with the objective function (1). In certain scenarios, the individual cost function $J_{i}(\cdot)$ depends not only on $\boldsymbol{x}_{i}$ but also on the collective decisions of all prosumers $\boldsymbol{x}$. For example, the wholesale energy purchasing price may be influenced by the cumulative power import/ export of all prosumers. It makes the problem a Generalized Nash Equilibrium (GNE) seeking problem [27]. Ensuring convergence in GNE seeking requires more stringent assumptions about problem (P1) and, in some cases, necessitates the pre-determination of P2P trading prices [28].
III Connection-aware P2P Trading and Peer Selection
@@
III-A Edge-based connection-aware P2P Trading
@@ 4       Step 3: Each prosumer receives new proposals and updates dual prices if are selected:
Algorithm 2 represents a significant improvement over the standard P2P trading algorithm. Its primary improvement lies in its relaxed requirement for updating by all trading peers. Specifically, in Step 2, a component referred to as ’a sensory oracle’ plays a pivotal role. This oracle discerns which trading connections $(i,j)\in\mathcal{E}^{(k)}$ should be activated in any given iteration $k$, thereby guiding the proposal updates accordingly. As an example, a simple implementation of this oracle could be an external, time-variant communication network dictating the feasible connections at any given moment. In Step 3, only the peers selected by the oracle are required to update their P2P trading prices. This approach introduces a form of semi-synchronization, contrasting with the Algorithm 1, which requires updates from all trading peers in every round. However, it maintains the principle that every message sent from the consumer $i$ to $j$ needs a corresponding response from $j$ to $i$, ensuring a closed communication loop.
@@
III-B Node-based Connection-Aware P2P Trading
@@ 4       Step 3: Each prosumer waits, receives new proposals, and records them as:
The node-based connection-aware P2P trading algorithm is shown in Algorithm 3. It requires much less synchrony and grants prosumers more freedom, which is more realistic in real trading situations.
@@
III-C Smart Trading Peer Selection Strategy
@@
The essence of this smart strategy lies in prioritizing peers with whom there are significant differences in trading proposals. These peers are more likely to influence the P2P trading price and, consequently, the cost for prosumer $i$. By focusing negotiation efforts on these potentially more impactful peers, prosumer $i$ can effectively accelerate the trading process.
@@ IV Convergence Analysis
IV-A Edge-based Algorithm and its Smart Selection Strategy
@@
IV-B Node-based Algorithm and its Smart Selection Strategy
@@
V Case Studies
V-A Basic Settings and Optimal Trading Results
@@
Some paramerters are set as follows: $\varrho_{i,t}$ is randomly selected between $[10,20]$/kW. $\xi_{i,t}$ is set to $-\varrho_{i,t}/2\overline{p_{i,t}^{\mathrm{L}}}$/(kW)${}^{2}$. $\overline{p_{i,t}^{\mathrm{L}}}$ is set to be 3 times of the recorded load profile; $\underline{p_{i,t}^{\mathrm{L}}}$ is set to be half of the recorded load profile. The maximum capacity of energy storage is set to 4 times of the average daily load in recorded profiles. The SOC is limited between $[0.1,1]$ of its capacity. $s_{i,0}=s_{i,\bar{\tau}}$ is set to $0.55$ times of the storage capacity. The charging and discharging efficiencies are both set at $95\%$. For the primal-dual algorithm, the learning rate $\rho=0.5$. All initial values of variables are set to 0. All the experiments are implemented using Matlab software’s parallel workers on the servers with an AMD EPYC 7H12 @ 2.60GHz CPU and 384.0 GB of RAM at the Beijing Super Cloud Computing Center. The optimization solver is the Gurobi software.
@@
V-B Algorithms and Smart Selection Strategies

### v2

# Cheng Feng Kedi Zheng Lanqing Shan Hani Alers Qixin Chen Lampros Stergioulas Hongye Guo

@@ Abstract
keywords:
P2P Energy Trading , Distributed Energy Resource , Trading Connection , Self-Selection Rights , Communication Asynchrony
[…]
1 Introduction
@@
Recent research highlights the importance of P2P trading in future distribution networks. Part of this research focused on a more detailed modeling of prosumers, considering their strategic behaviors [8], bounded rationality [9], and used learning techniques to help prosumers to perform trades [10]. Others incorporated a more detailed network model into trading problems, employing sensitivity factors [11] and DistFlow [12]. These models are valuable references for P2P trading analysis. Meanwhile, the implementation and development of solutions to trading models is also of significant importance. Various information exchange structures are designed to facilitate P2P trading, including leader-follower [13], bilateral [14], multi-layer [15], and grand coalition schemes [16]. This paper adopts a decentralized structure for its privacy protection and independence from a coordination entity [17]. Notably, most of the trading algorithms mentioned above, regardless of the information structure, neglect connection limits and require strict synchronization between trading peers or between the prosumer and the coordinator.
Addressing connection issues in P2P trading is critical. One type of related research focused on communication connection formation problems in prosumer coalitions. Some studies [18] illustrated that the trading connections among prosumers are flexible and can be adjusted according to prosumers’ preferences. One study proposes a graph coalition formation game [18], while another suggests peer matching [19]. These algorithms are mainly designed for P2P trading in a small community. They can be hard to be extended to handle large community because of their iterative complexity.
@@
Meanwhile, the connection issue is also a problem for non-coalition P2P trading. A few studies have investigated the non-coalition P2P trading connection problem. Ref.[20] offered an overview of information-related issues in P2P trading, and Ref.[21] conducted extensive numerical experiments on this topic. Ref.[22, 23] pioneered the use of asynchronous alternating direction method of multipliers (ADMM) to counter random communication issues in P2P trading. Nevertheless, Ref.[22] lacked a theoretical convergence guarantee for the constrained P2P trading problem. Ref.[23] introduced complex consensus processes during trading, exacerbating connection issues. Some works [24, 25] adopted an even-triggered version of the ADMM algorithm to address connection issues, named Communication-Sensored Consensus ADMM (COCA) [26]. In COCA, the connection peers are activated based on a diminishing threshold. Like all event-trigger algorithms, determining this threshold in real-world scenarios is challenging. Besides, the diminishing threshold should converge to zero, potentially activating almost all links over time. Ref.[27] used the node coloring algorithm to design a communication-efficient P2P trading mechanism. However, these works all assumed random connection problems. In contrast, the prosumers are not just nodes in communication networks, they have inherently different utilities and preferences. Existing algorithms ignored prosumers’ willingness and overlooked their self-selection rights. A complete of the existing research about communication issues in P2P trading is shown in Table 1.
Table 1: A comparison of P2P trading solutions against communication issues across different algorithms.
Handle
Asynchrony
Handle
Self-selection
Handle
Large Community
Handle
Long-time Negotiation
Theoretical
Guarantee
Coalition Game Graph Formation[18] $\times$ $\times$ $\times$
Peer Matching[19] $\times$ $\times$
Competitive Equilbrium Random-trigger[21, 22] $\times$ $\times$
Threshold-trigger[24, 25] $\times$ $\times$
Extra Consensus[23] $\times$ $\times$ $\times$
Node Coloring[27] $\times$
Our Approach
@@
The remainder of this paper is organized as follows: Section 2 outlines the P2P trading framework, formulates the problem of computing the P2P trading equilibrium, and discusses the standard primal-dual solution algorithm. Section 3 introduces two variants of the asynchronous P2P trading algorithm - the edge-based and node-based algorithms - and elaborates on the smart selection strategies. Section 4 presents a convergence proof for both algorithms and demonstrates the efficacy of the smart selection strategies in accelerating convergence. Section 5 conducts case studies validating the effectiveness of the proposed methods. Section 6 offers conclusions and suggests avenues for future research.
@@
2 Standard P2P Trading Model and Algorithm
2.1 Trading Setting and Competitive Equilibrium Computation
The system comprises prosumers, denoted as $i=1,...,I$. Each prosumer $i$ possesses DERs, including load equipment, rooftop photovoltaic panels, electric vehicles, and energy storage systems. These are integrated with an energy management system to facilitate monitoring, P2P trading, and billing. In the context of P2P trading, ponsumers $i$ are represented as nodes $\mathcal{N}$, and potential trading connections are symbolized by $\mathcal{E}$, following the conventions of graph theory. This study considers the situation where P2P trading is carried out in the day-ahead market, where $\tau=1,...,\bar{\tau}$ periods are considered. For prosumer $i$, its individual decision variables can include load usage power $\boldsymbol{p}_{i}^{\mathrm{L}}=\mathrm{col}(p_{i,\tau}^{\mathrm{L}})$, storage/vehicle charging power $\boldsymbol{p}_{i}^{\mathrm{CH}}=\mathrm{col}(p_{i,\tau}^{\mathrm{CH}})$, storage/vehicle discharging power $\boldsymbol{p}_{i}^{\mathrm{DIS}}=\mathrm{col}(p_{i,\tau}^{\mathrm{DIS}})$, storage/vehicle state of charge $\boldsymbol{s}_{i}=\mathrm{col}(s_{i,\tau})$, and exchanging power with the main grid $\boldsymbol{p}_{i}^{\mathrm{EX}}=\mathrm{col}(p_{i,\tau}^{\mathrm{EX}})$. For brevity, they are denoted in the concatenated form $\boldsymbol{x}_{i}$. $\boldsymbol{x}_{i}=\mathrm{col}(\boldsymbol{p}_{i}^{\mathrm{L}},\boldsymbol{p}_{i}^{\mathrm{CH}},\boldsymbol{p}_{i}^{\mathrm{DIS}},\boldsymbol{s}_{i},\boldsymbol{p}_{i}^{\mathrm{EX}})$. The information is private to prosumer $i$. P2P trading transactions between prosumer $i$ and $j$ are denoted as $\boldsymbol{t}_{i,j}=\mathrm{col}({t}_{i,j,\tau})$ for prosumer $i$ and $\boldsymbol{t}_{j,i}=\mathrm{col}({t}_{j,i,\tau})$ for prosumer $j$, which should only be known by $i$ and $j$. All possible P2P trading transactions of prosumer $i$ are denoted as $\boldsymbol{t}_{i}=\mathrm{col}(\boldsymbol{t}_{i,j})_{j\in\mathcal{N}_{i}}$. A pair refers to two prosumers, $i$ and $j$, who have potential trading intentions. There may be non-zero electricity trading between them, along with the possibility of P2P communication links.
This paper primarily focuses on fully decentralized P2P trading negotiations among prosumers. In real-world P2P trading systems, the communication topology may be hybrid, where groups of prosumers communicate with a centralized coordinator, while individual members within each group communicate in a P2P manner. In this scenario, the proposed method can be employed to assist with the communication process within each prosumer group.
Prosumers typically have the ability to engage in strategic behavior and exercise bargaining power in markets. However, in P2P trading, where there are numerous buyers and sellers dealing in identical electricity products, no individual prosumer can influence the market price. As a result, the outcome converges to a competitive equilibrium [28], which is the optimal solution to the social cost minimization problem, represented as:
@@ 2.1 Trading Setting and Competitive Equilibrium Computation
where the objective function (1) is to minimize the sum of individual costs $J_{i}\left(\boldsymbol{x}_{i},\boldsymbol{t}_{i}\right)$, which further depends on $i$’s decision variables $\boldsymbol{x}_{i}$ and P2P transactions $\boldsymbol{t}_{i}$. The cost typically includes negative load usage utility, storage age cost, network fee, tax and preferences for P2P trading, etc. Constraint (2) is the operation constraint for prosumer $i$, denoted by $\Omega_{i}$. The constraints can include load shifting constraints, energy storage charging status constraints, and power output constraints, etc. All the related information belongs to prosumer $i$’s privacy. Constraint (3) is the P2P trading power balance constraint: the power exported from $i$ to $j$ ($\boldsymbol{t}_{i,j}$) should be balanced by power from $j$ from $i$ ($\boldsymbol{t}_{j,i}$). The corresponding dual variable is $\boldsymbol{\lambda}_{i,j}$, represented by $\boldsymbol{\lambda}_{i,j}=\mathrm{col}(\lambda_{i,j,\tau})$ in different trading periods. The algorithms and solution methods can be used for day-ahead, real-time and rolling-based P2P trading situations.
@@
2.2 Primal-dual Algorithm for P2P trading
@@ 3       Step 2: Each prosumer receives new proposals and updates dual prices
By privacy protection, we mean for prosumer pair $i$, he or she will share trading transactions $\boldsymbol{t}_{i,j}$ for potential trading partners $j$. It represents how much electricity will be traded between $i$ and $j$ according to the willingness of prosumer $i$. Moreover, the primal-dual algorithm can be interpreted as prosumers engaging in a potential game, where the potential function aligns with the objective function (1). In certain scenarios, the individual cost function $J_{i}(\cdot)$ depends not only on $\boldsymbol{x}_{i}$ but also on the collective decisions of all prosumers $\boldsymbol{x}$. For example, the wholesale energy purchasing price may be influenced by the cumulative power import/ export of all prosumers. It makes the problem a Generalized Nash Equilibrium (GNE) seeking problem [29]. Ensuring convergence in GNE seeking requires more stringent assumptions about problem (P1) and, in some cases, necessitates the pre-determination of P2P trading prices [30].
3 Connection-aware P2P Trading and Peer Selection
@@
3.1 Edge-based connection-aware P2P Trading
@@ 4       Step 3: Each prosumer receives new proposals and updates dual prices if are selected:
Figure 2: (a) In Edge-based Algorithm, the sensory oracle selects which bilateral connections to be activated. (b) In Node-based Algorithm, each prosumer decides which potential trading partners to communicate with.
Algorithm 2 represents a significant improvement over the standard P2P trading algorithm. A graphical illustration about the algorithm are illustrated in Fig.2 (a). The key improvement of the algorithm lies in its relaxed requirement for updates from all trading peers. Specifically, in Step 2, a component referred to as the ”imaginary communication coordinator” plays a crucial role. This coordinator is responsible for monitoring and controlling the P2P communication links, determining which trading connections $(i,j)\in\mathcal{E}^{(k)}$ should be activated during a given iteration $k$, thereby guiding the P2P proposal updates accordingly. For instance, the coordinator might use a random selection policy to decide which P2P communication link will be activated. This scenario reflects the real, time-varying communication network that dictates the feasible connections at any given moment.
In Step 3, only the peers selected by the oracle are required to update their P2P trading prices. This approach introduces a form of semi-synchronization, contrasting with the Algorithm 1, which requires updates from all trading peers in every round. However, it maintains the principle that every message sent from the consumer $i$ to $j$ needs a corresponding response from $j$ to $i$, ensuring a closed communication loop.
@@
3.2 Node-based Connection-Aware P2P Trading
@@ 4       Step 3: Each prosumer waits, receives new proposals, and records them as:
The node-based connection-aware P2P trading algorithm is shown in Algorithm 3. It requires much less synchrony and grants prosumers more freedom, which is more realistic in real trading situations. A graphical illustration about the algorithm are illustrated in Fig.2 (b).
@@
Table 2 compares the communication details of the original P2P trading algorithm with the proposed edge-based and node-based algorithms. Among these, the node-based algorithm is the most promising for P2P trading. It eliminates the need for bi-directional synchronization and allows prosumers to choose their communication links based on their preferences to trading partners.
Table 2: A comparison of communication details across different algorithms.
Shared Data
Communication
Links
Bi-directional
or Uni-directional
Communication Link
Selection Method
Original Algorithm P2P transaction data All links Bi-directional No selection
Edge-based Algorithm Partial links Bi-directional
By a imaginary
communication coordinator
Node-based Algorithm Partial links Uni-directional By prosumers themselves
In real-world scenarios, communication latency between prosumer pairs can fluctuate over time. Here, we analyze the impact of latency in practical implementations.
The original trading algorithm requires strict synchronization among prosumers. Each prosumer $i$ must first send messages to its neighbors (incurring sending latency) and then wait to receive information from every neighbor (incurring receiving latency). In large communities, where each prosumer may have numerous trading partners, the likelihood of encountering abnormal communication latency among neighbours increases. As a result, the total sending and receiving latency can become quite significant.
In the edge-based algorithm, the number of prosumers required to send messages is reduced, thereby decreasing the probability of encountering extreme communication latency. However, waiting deadlocks can still occur.
In the node-based algorithm, prosumers are no longer required to wait indefinitely for their neighbors’ information. They can set a waiting deadline (e.g., 5 seconds), after which any incoming information is disregarded. This approach further reduces latency during the waiting phase.
3.3 Smart Trading Peer Selection Strategy
@@
The economical intuitions behind this smart strategy lie in prioritizing peers with whom there are significant differences in trading proposals. These peers are more likely to influence the P2P trading price and, consequently, the cost for prosumer $i$. By focusing negotiation efforts on these potentially more impactful peers, prosumer $i$ can effectively accelerate the trading process.
The two smart selection strategies presented above only requires a top-$e$ sorting process, whose complexity is far less than the solution process of prosumers’ individual optimization problem. Thereby, it will not add extra complexity to the asychronous P2P trading process.
@@ 4 Convergence Analysis
4.1 Edge-based Algorithm and its Smart Selection Strategy
@@
4.2 Node-based Algorithm and its Smart Selection Strategy
@@
5 Case Studies
5.1 Basic Settings and Optimal Trading Results
The proposed method can be applied to use cases across various time horizons, ranging from day-ahead (24 hours) to real-time (15 minutes). The case study will use a 24-hour profiled P2P trading example to demonstrate that the proposed method can handle more complex scenarios with longer optimization horizons.
@@ 5.1 Basic Settings and Optimal Trading Results
Some paramerters are set as follows: $\varrho_{i,t}$ is randomly selected between $[10,20]$/kW. $\xi_{i,t}$ is set to $-\varrho_{i,t}/2\overline{p_{i,t}^{\mathrm{L}}}$/(kW)2. $\overline{p_{i,t}^{\mathrm{L}}}$ is set to be 3 times of the recorded load profile; $\underline{p_{i,t}^{\mathrm{L}}}$ is set to be half of the recorded load profile. The maximum capacity of energy storage is set to 4 times of the average daily load in recorded profiles. The SOC is limited between $[0.1,1]$ of its capacity. $s_{i,0}=s_{i,\bar{\tau}}$ is set to $0.55$ times of the storage capacity. The charging and discharging efficiencies are both set at $95\%$. For the primal-dual algorithm, the learning rate $\rho=0.5$. All initial values of variables are set to 0. All the experiments are implemented using Matlab software’s parallel workers on the servers with an AMD EPYC 7H12 @ 2.60GHz CPU and 384.0 GB of RAM at the Beijing Super Cloud Computing Center. The optimization solver is the Gurobi software.
@@
5.2 Algorithms and Smart Selection Strategies